# Relative Overfitting and Accept-Reject Framework

### v1

@@ 6 Discussion and Future work
A Proof of Theorem 1
Let $V$ be the vocabulary size. We define $w_{1},w_{2},\dots,w_{V}\in V$ as word vectors sorted in descending order according to the true distribution $p(w|c)$, where $c$ represents the context set for prediction, $c\in C$. LLM and SLM refer to models based on similar architectures but differing in complexity.
The vocabulary $W$ is partitioned into three disjoint regions based on word frequency ranks derived from the true distribution $p(w|c)$:
  • High-frequency region (Core): $W_{core}=\{w_{1},\dots,w_{i}\}$(Index 1 to i as per definition)*
  • Mid-frequency region: $W_{mid}=\{w_{i+1},\dots,w_{j}\}$
  • Low-frequency region: $W_{low}=\{w_{j+1},\dots,w_{V}\}$
where $1\leq i<j<V$. The indices $i$ and $j$ define the boundaries based on observed model capabilities and frequency distinctions.
Based on existing research findings, we propose the following reasonable hypotheses:
  • Hypothesis 1:(Zipf’s Law) The distribution of words follows Zipf’s law.
  • Hypothesis 2:(Fitting Trend and Core Consistency) (a) (Error Trend) During language model training, high-frequency words are preferentially fitted. The absolute fitting error $|p(w_{k}|c)-q(w_{k}|c)|$ generally increases as rank $k$ increases.
    This implies that the average fitting error in the mid-frequency region $W_{mid}$ is higher than in the core region $W_{core}$.
    Let $E_{core}^{M}=\sum_{a=1}^{i}|p(w_{a}|c)-q_{M}(w_{a}|c)|$ denote the total core error, and $E_{mid}^{M}=\sum_{b=i+1}^{j}|p(w_{b}|c)-q_{M}(w_{b}|c)|$ denote the total mid-region error. Define the average errors (for $j>i$) as:
    $\bar{e}_{core}^{M}=\frac{E_{core}^{M}}{i}\quad\text{and}\quad\bar{e}_{mid}^{M}=\frac{E_{mid}^{M}}{j-i}.$ (7)
    The implication of the error trend is $\bar{e}_{mid}^{LLM}>\bar{e}_{core}^{LLM}$.
    (b) (Core Consistency) The fitting capabilities of LLM and SLM in the high-frequency core region $W_{core}$ (ranks 1 to i) are consistent, meaning their average errors are approximately equal:
    $|\bar{e}_{core}^{LLM}-\bar{e}_{core}^{SLM}|<\epsilon$ (8)
    where $\epsilon$ is a small positive constant.
  • Hypothesis 3:(Model Output Behavior)
    (1) Both models predict effectively in $W_{core}$ ($k\in[1,i]$). *(Redundant given H2(b), kept for clarity of structure)*
    (2) In $W_{mid}$ ($k\in[i+1,j]$), $q_{LLM}(w_{k}|c)>0$ for all $k$, while $q_{SLM}(w_{k}|c)\approx 0$.
    (3) In $W_{low}$ ($k\in[j+1,V]$), $q_{LLM}(w_{k}|c)\approx 0$ and $q_{SLM}(w_{k}|c)\approx 0$.
Based on the above hypotheses, we define the core as the high-frequency region $w_{1}$-$w_{i}$ and obtain the following theoretical result regarding the model’s average fitting error:
Let $a\in[1,i]$, $b\in[i+1,j]$. The average fitting error of LLM over the range [1, j] is greater than or equal to that of SLM over the range [1, i]:
$\frac{\sum_{a=1}^{i}|p(w_{a}|c)-q_{LLM}(w_{a}|c)|+\sum_{b=i+1}^{j}|p(w_{b}|c)-q_{LLM}(w_{b}|c)|}{j}\geq\frac{\sum_{a=1}^{i}|p(w_{a}|c)-q_{SLM}(w_{a}|c)|}{i}$ (9)
Let $LHS=\frac{\sum_{a=1}^{i}|p(w_{a}|c)-q_{LLM}(w_{a}|c)|+\sum_{b=i+1}^{j}|p(w_{b}|c)-q_{LLM}(w_{b}|c)|}{j}$. Let $RHS=\frac{\sum_{a=1}^{i}|p(w_{a}|c)-q_{SLM}(w_{a}|c)|}{i}$. We aim to prove $LHS\geq RHS$.
Define the average errors as in Hypothesis 2: $\bar{e}_{core}^{M}=\frac{\sum_{a=1}^{i}|p(w_{a}|c)-q_{M}(w_{a}|c)|}{i}$ for the core region [1, i]. $\bar{e}_{mid}^{LLM}=\frac{\sum_{b=i+1}^{j}|p(w_{b}|c)-q_{LLM}(w_{b}|c)|}{j-i}$ for the mid region [i+1, j] (assuming j>i).
From Hypothesis 2(b) (Core Consistency), we have:
$|\bar{e}_{core}^{LLM}-\bar{e}_{core}^{SLM}|<\epsilon$ (10)
So, $RHS=\bar{e}_{core}^{SLM}\approx\bar{e}_{core}^{LLM}$.
From Hypothesis 2(a) (Error Trend), we have:
$\bar{e}_{mid}^{LLM}>\bar{e}_{core}^{LLM}\quad(2)$ (11)
Now, express LHS in terms of average errors:
$LHS=\frac{i\cdot\bar{e}_{core}^{LLM}+(j-i)\bar{e}_{mid}^{LLM}}{j}$ (12)
Substitute $\bar{e}_{mid}^{LLM}$ using inequality (2). Since $\bar{e}_{mid}^{LLM}>\bar{e}_{core}^{LLM}$, we replace $\bar{e}_{mid}^{LLM}$ with the smaller value $\bar{e}_{core}^{LLM}$ to obtain a strict inequality:
$LHS>\frac{i\cdot\bar{e}_{core}^{LLM}+(j-i)\bar{e}_{core}^{LLM}}{j}=\frac{(i+j-i)\bar{e}_{core}^{LLM}}{j}=\frac{j\bar{e}_{core}^{LLM}}{j}=\bar{e}_{core}^{LLM}$ (13)
So, we have shown:
$LHS>\bar{e}_{core}^{LLM}\quad(3)$ (14)
Combining (1) and (3):
$LHS>\bar{e}_{core}^{LLM}\approx\bar{e}_{core}^{SLM}=RHS$ (15)
[…]

### v4

@@ 6 Discussion and Future work
A Proof of Theorem
A.1 Definitions
Definition A.1.1.
Let $q$ and $q^{\prime}$ be two models sharing the same input and output spaces. If model $q$ has a larger number of parameters (or higher model complexity) than model $q^{\prime}$, we define $q$ as the Large Model (LM) and $q^{\prime}$ as the Small Model (SM).
Definition A.1.2.
Let $S$ be the set of all samples (or positions) to be predicted in a given task, its size being $N=|S|$. Let $s\in S$ represent a specific sample. Let $V=\{1,2,3,\dots,|V|\}$ be the set of indices for all possible output classes (dimensions), where the output dimensions of the two models correspond one-to-one.
Definition A.1.3.
Let $k_{s}\in V$ be the index of the true class for sample $s$. Let $q_{s}$ be the predicted output vector of model $q$ for sample $s$.
Definition A.1.4.
Let $q_{s,i}$ be the probability assigned by model $q$ to sample $s$ belonging to class $i$. Similarly, $q^{\prime}_{s,i}$ is defined for model $q^{\prime}$.
Definition A.1.5.
Model $q$ predicts sample $s$ correctly if and only if $\arg\max_{i\in V}(q_{s,i})=k_{s}$.
Definition A.1.6.
The entire set of samples in the task is partitioned into three sets: $T$, $F$, and $N$.
  • The set $T$ consists of samples where the LM predicts incorrectly, but the SM predicts correctly:
    $T=\{s\in S\mid\arg\max_{i\in V}(q_{s,i})\neq k_{s}\text{ and }\arg\max_{i\in V}(q^{\prime}_{s,i})=k_{s}\}$
  • The set $F$ consists of samples where the LM predicts correctly, but the SM predicts incorrectly:
    $F=\{s\in S\mid\arg\max_{i\in V}(q_{s,i})=k_{s}\text{ and }\arg\max_{i\in V}(q^{\prime}_{s,i})\neq k_{s}\}$
  • The set $N$ consists of samples where both LM and SM predict correctly, or both predict incorrectly.
We have $T\cup F\cup N=S$, and $T,F,N$ are mutually disjoint.
Definition A.1.7.
The accuracy (ACC) of a model on a prediction task $S$ is defined as (where $I(\cdot)$ is the indicator function):
$ACC_{q}=\frac{1}{|S|}\sum_{s\in S}I(\arg\max_{i\in V}(q_{s,i})=k_{s})$
A.2 Assumptions and Core Theorems
Assumption A.2.1.
The improvements of larger models are primarily concentrated in the tail of the distribution. Assume there exists a high-frequency output class $h\in V$. Let $S_{h}^{q}=\{s\in S\mid\arg\max_{i}q_{s,i}=h\}$ be the set of samples predicted as $h$ by LM, with size $n=|S_{h}^{q}|$. Let $S_{h}^{q^{\prime}}=\{s\in S\mid\arg\max_{i}q^{\prime}_{s,i}=h\}$ be the set of samples predicted as $h$ by SM, with size $m=|S_{h}^{q^{\prime}}|$.
We assume the following conditions hold53 ; 54 ; 55 ; 56 ; 57 ; 58 :
  1. 1.
    The SM outputs class $h$ more frequently than the LM, i.e., $n<m$.
  2. 2.
    The advantage of LM’s prediction accuracy for class $h$, $P_{L}(h)=P(k_{s}=h\mid s\in S_{h}^{q})$, over SM’s accuracy, $P_{S}(h)=P(k_{s}=h\mid s\in S_{h}^{q^{\prime}})$, is limited. Specifically, this advantage satisfies the relation:
    $|P_{L}(h)-P_{S}(h)|<\frac{(m-n)P_{S}(h)}{n}$
Theorem A.2.2 (Mainstay Deviation).
Under Assumption A.2.1, for class $h$, the total number of correct predictions by SM is greater than that by LM.
Proof.
The total number of correct predictions for class $h$ by LM is $N_{L}^{h}=n\cdot P_{L}(h)$. The total number of correct predictions for class $h$ by SM is $N_{S}^{h}=m\cdot P_{S}(h)$.
We want to prove $N_{L}^{h}<N_{S}^{h}$, which is equivalent to proving $m\cdot P_{S}(h)-n\cdot P_{L}(h)>0$.
We transform the difference:
$m\cdot P_{S}(h)-n\cdot P_{L}(h)$ $=m\cdot P_{S}(h)-n\cdot P_{S}(h)+n\cdot P_{S}(h)-n\cdot P_{L}(h)$
$=(m-n)P_{S}(h)-n(P_{L}(h)-P_{S}(h))$
From Assumption A.2.1, we have:
$(P_{L}(h)-P_{S}(h))\leq|P_{L}(h)-P_{S}(h)|<\frac{(m-n)P_{S}(h)}{n}$
Thus,
$(m-n)P_{S}(h)-n(P_{L}(h)-P_{S}(h))>(m-n)P_{S}(h)-n\left(\frac{(m-n)P_{S}(h)}{n}\right)=0$
The inequality is proven.
It can be easily shown that Theorem A.2.2 holds for any high-frequency output dimension satisfying Assumption A.2.1. ∎
Lemma A.2.3 (Exchange Condition).
Let the fused model be $q_{new,s,i}=w\cdot q_{s,i}+(1-w)\cdot q^{\prime}_{s,i}$ for $w\in(0,1)$. On sample $s$, if the LM’s original prediction is $i$ (i.e., $\arg\max(q_{s})=i$), for the fused model’s prediction for class $j$ to have a higher probability than for class $i$, it must satisfy:
$w\cdot q_{s,j}+(1-w)\cdot q^{\prime}_{s,j}>w\cdot q_{s,i}+(1-w)\cdot q^{\prime}_{s,i}$
For $w\in(0,1)$, this inequality is equivalent to:
$\frac{q_{s,i}-q_{s,j}}{q^{\prime}_{s,j}-q^{\prime}_{s,i}}<\frac{1-w}{w}$
Clearly, we need SM’s prediction for $j$ to be greater than its prediction for $i$:
$q^{\prime}_{s,j}-q^{\prime}_{s,i}>0$
This is termed the exchange condition.
Furthermore, if $\arg\max(q_{new,s})=j$, then the fused model’s prediction changes from $i$ to $j$. This is termed the strict exchange condition if the exchange condition is also met.
We define the left side as the Exchange Threshold, denoted $ET(s,i\to j)$. A flip occurs only when $ET<\frac{1-w}{w}$.
This lemma implies that only when the exchange condition is met can the fused model’s prediction potentially differ from the LM’s prediction. Any case not satisfying this condition cannot change the prediction outcome. Moreover, only when the strict exchange condition is met can the change in prediction outcome be determined.
Assumption A.2.4 (Stratification of Predictions).
Typically, a model exhibits greater discrimination for samples it predicts correctly68 ; 69 ; 70 . Specifically, we assume situations exist where: for some samples correctly predicted by model $q$ (e.g., $b\in F$), the difference between the probability of the correctly predicted class $q_{b,k_{b}}$ and the probability of the second most likely class (or any other incorrect class $x_{b}$, $q_{b,x_{b}}$), i.e., $(q_{b,k_{b}}-q_{b,x_{b}})$, is greater than the difference for some samples incorrectly predicted by model $q$ (e.g., $a\in T$) between its incorrectly predicted class probability $q_{a,i_{a}}$ and the true class probability $q_{a,k_{a}}$ (or some other reference class $z_{a}$, $q_{a,z_{a}}$), i.e., $(q_{a,i_{a}}-q_{a,z_{a}})$.
Formally, there exists at least one sample $a\in T$ such that for all samples $b\in F$ (if $F$ is non-empty):
$(q_{b,k_{b}}-\max_{j\neq k_{b}}q_{b,j})>(q_{a,i_{a}}-q_{a,k_{a}})$
where $k_{b}=\arg\max(q_{b})$ is LM’s correct prediction for $b$, $i_{a}=\arg\max(q_{a})$ is LM’s incorrect prediction for $a$, and $k_{a}$ is the true class of $a$.
A similar condition holds for model $q^{\prime}$:
$(q^{\prime}_{b,k_{b}}-\max_{j\neq k_{b}}q^{\prime}_{b,j})>(q^{\prime}_{a,i^{\prime}_{a}}-q^{\prime}_{a,k_{a}})$
Let $A$ be the set of all samples $a\in T$ satisfying this condition.
Theorem A.2.5 (AR Law: Improvement, Not Degradation).
We say Condition (1) holds if there exists a position $s\in S$ where SM’s predictive capability is stronger than LM’s. Specifically, LM predicts incorrectly, $k_{s}\neq\arg\max(q_{s})$, and SM’s predicted probability for the true class $k_{s}$ is greater than its predicted probability for LM’s incorrect prediction $i_{s}=\arg\max(q_{s})$, i.e., $q^{\prime}_{s,k_{s}}>q^{\prime}_{s,i_{s}}$. (From Theorem A.2.2 and experimental results, this condition is generally met.)
Under Assumption A.2.4: For a linear fusion model $q_{new}=w\cdot q+(1-w)\cdot q^{\prime}$ (where $w\in(0,1)$) of a large model $q$ and a small model $q^{\prime}$, if and only if Condition (1) holds, a weight $w$ can always be found such that the accuracy of the fused model $ACC_{new}$ is potentially higher than, and not worse than, the accuracy of model $q$, i.e., $ACC_{new}\geq ACC_{q}$.
If, additionally, for the set $A$, there exists at least one position satisfying the strict exchange condition, then the accuracy of the fused model $ACC_{new}$ is higher than the accuracy of model $q$, i.e., $ACC_{new}>ACC_{q}$. In fact, for every predicted position $s$, the conclusion is that the accuracy will not decrease.
Proof.
From Lemma A.2.3, for the weighted output to potentially differ, the exchange condition must be met. For $\arg\max(q_{s})=i$:
$w\cdot q_{s,j}+(1-w)\cdot q^{\prime}_{s,j}>w\cdot q_{s,i}+(1-w)\cdot q^{\prime}_{s,i}$
For output $j$ to replace output $i$ as the final output, it is also required that:
$\arg\max(w\cdot q_{s}+(1-w)\cdot q^{\prime}_{s})=j$
Assume Condition (1) does not hold. Then for any position $s\in S$: If $q_{s,i}=\arg\max(q_{s})=k_{s}$, then $q$ is already optimal at this position, and fusion can only lead to a worse result. If $q_{s,i}=\arg\max(q_{s})\neq k_{s}$, but $q^{\prime}_{s,k_{s}}<q^{\prime}_{s,i_{s}}$ (where $i_{s}$ is the LM’s incorrect prediction), then
$w\cdot q_{s,k_{s}}+(1-w)\cdot q^{\prime}_{s,k_{s}}<w\cdot q_{s,i_{s}}+(1-w)\cdot q^{\prime}_{s,i_{s}}$
This means the true class $k_{s}$ can never replace the LM’s incorrect prediction $i_{s}$ as the final output. Thus, the necessity of Condition (1) is proven.
Consider the partition of the task into sets $T,F,N$. Let $t\in T$ and $f\in F$. For set $N$: if both models predict correctly, the weighted result cannot change the original $\arg\max$. If both predict incorrectly, the result cannot worsen. Thus, for set $N$, weighting can only potentially improve ACC.
For set $T$: if the exchange condition $\frac{q_{t,i_{t}}-q_{t,k_{t}}}{q^{\prime}_{t,k_{t}}-q^{\prime}_{t,i_{t}}}<\frac{1-w}{w}$ (where $i_{t}$ is LM’s incorrect prediction and $k_{t}$ is the true class) is met, the incorrect result $i_{t}$ can be replaced by the correct result $k_{t}$.
For set $F$: if the exchange condition $\frac{q_{f,k_{f}}-q_{f,j_{f}}}{q^{\prime}_{f,j_{f}}-q^{\prime}_{f,k_{f}}}<\frac{1-w}{w}$ (where $k_{f}$ is LM’s correct prediction and $j_{f}$ is SM’s incorrect prediction) is met, the true result $k_{f}$ can be replaced by an incorrect result $j_{f}$, causing ACC to decrease.
By Assumption A.2.4, there exists some $a\in A\subseteq T$ such that for all $f\in F$:
$(q_{f,k_{f}}-\max_{j\neq k_{f}}q_{f,j})>(q_{a,i_{a}}-q_{a,k_{a}})$
and
$(q^{\prime}_{a,k_{a}}-q^{\prime}_{a,i^{\prime}_{a}})>(q^{\prime}_{f,j_{f}}-q^{\prime}_{f,k_{f}})$
This implies that the exchange threshold for correcting $a$ is smaller than the exchange threshold for incorrectly changing $f$:
$\frac{q_{a,i_{a}}-q_{a,k_{a}}}{q^{\prime}_{a,k_{a}}-q^{\prime}_{a,i_{a}}}<\frac{q_{f,k_{f}}-q_{f,j_{f}}}{q^{\prime}_{f,j_{f}}-q^{\prime}_{f,k_{f}}}$
At this point, we can always find a $w$ such that:
$\frac{q_{a,i_{a}}-q_{a,k_{a}}}{q^{\prime}_{a,k_{a}}-q^{\prime}_{a,i_{a}}}<\frac{1-w}{w}<\frac{q_{f,k_{f}}-q_{f,j_{f}}}{q^{\prime}_{f,j_{f}}-q^{\prime}_{f,k_{f}}}$
Since $a\in T$, this means that a potential correct substitution in set $T$ is satisfied, while an incorrect substitution in set $F$ is not. Thus, ACC may increase and cannot decrease.
If there is at least one point $a^{\prime}\in A$ that satisfies the strict exchange condition, then at this position, the incorrect result will be replaced by the correct result, causing an increase in ACC. There are N additional episodes that will not cause a decline in accuracy, thus there exists a w such that at any prediction position s, accuracy will not decrease. ∎
Theorem A.2.6 (AR Law: Maximum Improvement Magnitude).
Under the conditions of Theorem A.2.5, let $A$ be the set of all positions satisfying Assumption A.2.4. Let $R$ be the set of all positions satisfying the strict exchange condition, where $R\subseteq A\subseteq T\subseteq S$. The accuracy improvement $\Delta ACC=ACC_{new}-ACC_{q}$ of the fused model $q_{new}$ relative to the original large model $q$ has the following properties:
  • $\Delta ACC$ is a strictly monotonically increasing function of the size of set $R$, $|R|$.
  • $\Delta ACC$ is a monotonically increasing function of the sizes of sets $A$ and $T$, $|A|$ and $|T|$.
Proof.
According to Theorem A.2.5, by choosing an optimal fusion weight $w$, we can ensure that only samples in set $T$ are corrected, and samples in set $F$ are not incorrectly changed.
By definition, $R$ is the set of all samples that can be safely corrected within the optimal weight range. Then, the maximum accuracy improvement $\Delta ACC_{max}$ directly depends on the size of $R$:
$\Delta ACC_{max}=\frac{|R|}{|S|}$
From this equation, it is clear that if $|R|$ increases, $\Delta ACC_{max}$ will also increase proportionally. Therefore, $\Delta ACC_{max}$ is a strictly monotonically increasing function of $|R|$.
By definition, we have the subset relations $R\subseteq A\subseteq T$. This directly leads to the inequality relations for set sizes: $|R|\leq|A|\leq|T|$. Therefore, the maximum accuracy improvement is limited by $|A|$ and $|T|$:
$\Delta ACC_{max}=\frac{|R|}{|S|}\leq\frac{|A|}{|S|}\leq\frac{|T|}{|S|}$
This inequality shows that the upper limit of ACC improvement is determined by $|A|$ and $|T|$.
Now consider monotonicity. If set $A_{1}\subseteq A_{2}$, then the correctable sets extracted from them also satisfy $R(A_{1})\subseteq R(A_{2})$. Therefore, $\Delta ACC_{max}(A_{1})\leq\Delta ACC_{max}(A_{2})$. This means $\Delta ACC_{max}$ is a monotonically increasing (non-decreasing) function of $|A|$.
By the same logic, since $A\subseteq T$, $\Delta ACC_{max}$ is also a monotonically increasing (non-decreasing) function of $|T|$. ∎
Theorem A.2.7 (AR Law: ACC Improvement Variation with Weight).
Building on Theorem A.2.5, let $\beta=w/(1-w)$. As the weight $w$ decreases (and thus $\beta$ decreases, while $(1-w)/w$ increases), the ACC improvement magnitude initially increases monotonically.
If for any $r\in R$, Assumption A.2.4 is satisfied (implying $R=A=T$ under ideal conditions), then there exists an optimal weight range where $\beta=\alpha$ such that the ACC improvement reaches its maximum. Subsequently, as $w$ further decreases ( $\beta$ further decreases), the ACC improvement magnitude monotonically decreases, until $\beta=D$, after which ACC begins to decline (That is, the lifting domain of ACC is $[D,\infty]$).
Proof.
From Theorem A.2.5, we have the condition for favorable exchange:
$\frac{q_{a,i_{a}}-q_{a,k_{a}}}{q^{\prime}_{a,k_{a}}-q^{\prime}_{a,i_{a}}}<\frac{1-w}{w}<\frac{q_{f,k_{f}}-q_{f,j_{f}}}{q^{\prime}_{f,j_{f}}-q^{\prime}_{f,k_{f}}}$
Let $A^{\prime}$ be the set of points satisfying $\frac{q_{a,i_{a}}-q_{a,k_{a}}}{q^{\prime}_{a,k_{a}}-q^{\prime}_{a,i_{a}}}<\frac{1-w}{w}$ and also satisfying the strict exchange condition. It is clear that the current ACC improvement is proportional to $|A^{\prime}|$, and $A^{\prime}\subseteq R$.
With $A$ (and thus $R$) fixed, as $w$ decreases, $\beta=w/(1-w)$ decreases, and $\tau=(1-w)/w$ increases. As $\tau$ increases, more samples in $R$ will satisfy $ET<\tau$, so $|A^{\prime}|$ monotonically increases. Consequently, the ACC improvement magnitude increases.
If $R=A=T$ (ideal case where all potential corrections satisfy Assumption A.2.4), then when $|A^{\prime}|=|R|$ (all samples in $R$ are corrected), the ACC improvement reaches its maximum. At this point, let the corresponding $\beta$ be $\alpha$.
As $w$ continues to decrease ( $\beta$ continues to decrease, $\tau$ continues to increase), $|A^{\prime}|$ no longer increases (it is capped at $|R|$). Now, $R=T=A^{\prime}$. The increase in $\tau$ will only cause more samples in $F$ to satisfy
$\frac{q_{f,k_{f}}-q_{f,j_{f}}}{q^{\prime}_{f,j_{f}}-q^{\prime}_{f,k_{f}}}<\frac{1-w}{w}$
Let $B$ be the set of these positions from $F$. For $f\in B$, an incorrect prediction $j_{f}$ by SM replaces the correct LM prediction $k_{f}$, causing the ACC improvement magnitude to decrease (i.e., net ACC starts to fall).
When $|B|$ (number of newly incorrect predictions) becomes greater than $|R|$ (number of correct predictions that were maintained or newly made), the net change in ACC becomes negative, meaning ACC starts to decline rather than improve. We define the weight ratio $\beta$ at which this first occurs as $D$. ∎
A.3 Corollaries
Corollary A.3.1 (Multi-dimensional Extension).
The conclusions of the AR Law (Theorems A.2.5, A.2.6, A.2.7) hold for any dimension of the output vector. In other words, the validity of these laws does not depend on the total number of dimensions in the output space.
Proof.
We use mathematical induction on the total number of output dimensions $D_{dim}=|V|$.
Base Case: When $D_{dim}=2$, the problem is a binary classification. The LM’s prediction $i$ and SM’s prediction $k$ (if different) are the only two classes. All comparisons regarding Exchange Thresholds (ET) and prediction stratification (Assumption A.2.4) are made directly on these two dimensions. The logic and derivations of Theorems A.2.5, A.2.6, A.2.7 hold directly in this simplest case.
Inductive Hypothesis: Assume that for any output space of dimension $D_{dim}-1$ (where $D_{dim}>2$), the AR Law holds.
Inductive Step: We need to prove that under this assumption, the AR Law also holds for an output space of dimension $D_{dim}$.
Consider a $D_{dim}$-dimensional output space $V$. The core mechanism of the AR Law, i.e., the change in prediction of the fused model $q_{new}$, depends on the exchange condition in Lemma A.2.3. For example, for the prediction to flip from class $i$ to class $j$, it must satisfy:
$w\cdot q_{s,j}+(1-w)\cdot q^{\prime}_{s,j}>w\cdot q_{s,i}+(1-w)\cdot q^{\prime}_{s,i}$
Rearranging gives the condition for the exchange threshold:
$\frac{q_{s,i}-q_{s,j}}{q^{\prime}_{s,j}-q^{\prime}_{s,i}}<\frac{1-w}{w},\quad\text{where }q^{\prime}_{s,j}>q^{\prime}_{s,i}$
Crucially, this core condition for flipping only involves the probabilities of these two dimensions, $i$ and $j$, and is independent of the probability values of any other dimension $u\in V\setminus\{i,j\}$.
Now, let us arbitrarily remove one dimension $u$ from the $D_{dim}$-dimensional space, forming a $(D_{dim}-1)$-dimensional subspace $V^{\prime}=V\setminus\{u\}$. We can construct a new classification subproblem defined on $V^{\prime}$. For this subproblem, we can define a new set of probability distributions normalized over $V^{\prime}$ to determine which class to output.
Importantly, for any two dimensions still in $V^{\prime}$ (e.g., $i$ and $j$), their relative probability magnitudes in the original models $q$ and $q^{\prime}$, and the value of the exchange threshold $ET(s,i\to j)$ calculated from them, are independent of whether we consider dimension $u$.
Therefore, in this $(D_{dim}-1)$-dimensional subspace $V^{\prime}$, all prerequisites of the AR Law (such as Assumptions A.2.1 and A.2.4) and core mechanisms (such as Lemma A.2.3) apply equally. Hence, the AR Law holds for this $(D_{dim}-1)$-dimensional subproblem.
Since the removed dimension $u$ was chosen arbitrarily, this implies that the intrinsic logic of the AR Law holds on any subset of dimensions. Thus, the validity of the law is not restricted to a specific number of dimensions; it naturally extends from $D_{dim}-1$ dimensions to $D_{dim}$ dimensions.
By induction, the AR Law holds for any output space of dimension $D_{dim}\geq 2$.
Specifically, if the removed dimension is the output dimension of the fused model itself at that position, the output changes according to the AR Law. If not, the output at this position does not change. ∎
Corollary A.3.2 (Variant based on Assumption A.2.4).
If, in set $T$, the number of positions satisfying Assumption A.2.4 and the strict exchange condition (denoted as set $R$) is greater than the number of positions not satisfying Assumption A.2.4 (denoted as $T\setminus A$), i.e., $|R|>|T\setminus A|$, then Theorem A.2.5 and Theorem A.2.6 still hold.
Proof.
The goal of this proof is to find a fusion weight $w$ that achieves a net improvement in accuracy under the condition that Assumption A.2.4 does not fully hold, by utilizing the premise $|R|>|T\setminus A|$.
The core challenge is that "risky" samples $a^{\prime}\in T\setminus A$ might have an exchange threshold $ET(a^{\prime})$ greater than the exchange threshold $ET(f)$ of some samples $f\in F$, which means the threshold separation relied upon in the proof of Theorem A.2.5 is no longer absolutely guaranteed.
However, the premise $|R|>|T\setminus A|$ allows us to adopt a conservative strategy to mitigate this risk. We can construct a safe benign sample subset $R_{safe}\subseteq R$, with its size defined as $|R_{safe}|=|R|-|T\setminus A|$.
According to the premise, this subset contains at least one element. Since all elements of $R_{safe}$ are from $R$, they must satisfy Assumption A.2.4. Therefore, for any $r\in R_{safe}$ and $f\in F$, their exchange thresholds satisfy $ET(r)<ET(f)$. This property ensures that the exchange thresholds of all samples in $R_{safe}$ are strictly smaller than the exchange thresholds of all samples in $F$.
Thus, we can always choose a weight ratio $\tau=\frac{1-w}{w}$ such that it falls between the threshold ranges of these two sets, i.e., satisfying $\max_{r\in R_{safe}}\{ET(r)\}<\tau<\min_{f\in F}\{ET(f)\}$.
For a weight $w$ satisfying this condition, all samples in $R_{safe}$ will be successfully corrected, yielding an accuracy gain of at least $\frac{|R_{safe}|}{|S|}>0$, while ensuring that samples in set $F$ are not incorrectly changed, resulting in zero loss. Thus, the accuracy of the fused model achieves a net improvement, proving that Theorem A.2.5 still holds under this variant condition.
The proof for Theorem A.2.6 is then evident. ∎
Corollary A.3.3 (With SM as the Primary Model).
All the preceding theorems are proven with the LM as the primary model, which is generally due to LM having higher ACC than SM.
If the opposite situation occurs, where the small model $q^{\prime}$ has higher accuracy than the large model $q$, then by swapping the roles of the two models, and under the satisfaction of the assumptions, it can similarly be proven that there exists a fusion weight $w$ such that the accuracy of the fused model is higher than that of the original SM model $q^{\prime}$.
Proof.
The proof of this corollary is based on the principle of symmetry. When we consider the more accurate small model $q^{\prime}$ as the primary model and $q$ as the auxiliary model, we aim to prove that the accuracy of the fused model can exceed that of $q^{\prime}$.
To do this, we need to redefine the sets $T$ and $F$ that form the basis of the theory. The new set $T^{\prime}$, i.e., positions where the primary model ($q^{\prime}$) predicts incorrectly and the auxiliary model ($q$) predicts correctly, is defined identically to the original set $F$. Similarly, the new set $F^{\prime}$, i.e., positions where the primary model ($q^{\prime}$) predicts correctly and the auxiliary model ($q$) predicts incorrectly, is defined identically to the original set $T$.
Therefore, swapping roles in the proof process is equivalent to systematically interchanging $T\leftrightarrow F$ in the original proofs. Considering that the mathematical structure of the linear fusion model $q_{new}$ and the calculation of its exchange threshold are inherently symmetrical with respect to $q$ and $q^{\prime}$, we can conclude that by performing the aforementioned symmetrical substitutions of roles, sets, and assumptions, the proof chains of all theorems remain unchanged. ∎
Corollary A.3.4 (Extension of AR Law to Multi-Model Ensembles).
Assume that every pairwise combination of multiple models satisfies Assumption A.2.1. If the initial weights of each model form a geometric progression, the principles of the two-model ensemble theorem can be naturally extended to multi-model ensembles.
Proof.
Assume that there are $m$ models participating in the ensemble and these models are sorted by performance from lowest to highest. The exchange condition then becomes:
$\sum_{h=1}^{m}w_{h}\cdot q_{h,s,j}>\sum_{h=1}^{m}w_{h}\cdot q_{h,s,i}$
Here, $q_{h,s,j}$ denotes the output of the $h$-th model, and we have $ACC_{q_{x_{2}}}>ACC_{q_{x_{1}}}$ for any $x_{2}>x_{1}$.
This expression can be decomposed into $C_{m}^{2}$ pairwise model combinations:
$w_{x_{2}}\cdot q_{x_{2},s,j}+w_{x_{1}}\cdot q_{x_{1},s,j}>w_{x_{2}}\cdot q_{x_{2},s,i}+w_{x_{1}}\cdot q_{x_{1},s,i}$
For each pairwise combination, as established previously, the AR Law holds. This means there exists a "safe window" for the weight ratio, defined by a lower and upper bound:
$\frac{q_{x_{2},a,i_{a}}-q_{x_{2},a,k_{a}}}{q_{x_{1},a,k_{a}}-q_{x_{1},a,i_{a}}}<\frac{w_{x_{1}}}{w_{x_{2}}}<\frac{q_{x_{2},f,k_{f}}-q_{x_{2},f,j_{f}}}{q_{x_{1},f,j_{f}}-q_{x_{1},f,k_{f}}}$
Let’s denote the weight ratio $\frac{w_{x_{1}}}{w_{x_{2}}}$ as $k(x)$, and the lower and upper bounds as $LL(x)$ and $UL(x)$ respectively, where $x$ is the index for the combination, ranging from $1$ to $C_{m}^{2}$.
It is evident that if the pairwise exchange condition holds for all combinations, the overall condition must also hold (though the reverse is not necessarily true). Continuing with the logic of the previous proof, we can control the ratio $k(x)$. When $k(x)<\min(UL(x))$, as its value increases, the number of possible correct substitutions for samples in set T will increase, while incorrect substitutions for samples in set F will not occur.
Since the model weights form a geometric progression, we can set $k(x)\in[r,r^{m-1}]$, where $r$ is a constant common ratio. This results in the weights $\{w_{h}\}$ forming an increasing geometric sequence. Given the constraint $\sum_{h=1}^{m}w_{h}=1$, we can solve the base weight:
$w_{1}=\frac{1-r}{1-r^{m}}$
This initial weight distribution relates $k(x)$ to $r$. Therefore, we can control $k(x)$ for all pairwise combinations by adjusting $r$. The role of $r$ here becomes equivalent to the weight ratio $\frac{1-w}{w}$ in the original two-model proof.
Consequently, we only need to set the upper bound as $\min(UL(x))$. As long as we choose an $r$ such that:
$LL(x)<r<\min(UL(x))$
we can similarly prove analogous theorems for the multi-model ensemble case.
For the case of a strict increase in accuracy, we also need to choose $r$ such that $\max(LL(x))<r<\min(UL(x))$ to ensure that the overall exchange condition is met definitively.
A.4 Concluding Remarks
Theorems A.2.5, A.2.6, and A.2.7 are collectively termed the AR Law. It reveals the principles governing how the accuracy of a fused model changes under certain assumptions. We have observed in tens of thousands of experiments that these laws hold, and combined with some existing research findings, we believe the assumptions of the AR Law are generally valid.
[…]